OpenStudy (anonymous):

Help is very appreciated! Y is a differentiable function of x. Choose the alternative that is the derivative dy/dx. x^3-y^3=1 x+cos(x+y)=0 3x^2-2xy+5y^2=1 I'm doing these for homework on my AP Calc preparation course and am really confused. The teacher kind of rushed over this and I don't really understand what they meant. Thank You!

OpenStudy (anonymous):

@ganeshie8 @myininaya @pooja195 @nincompoop @Preetha @phi

OpenStudy (anonymous):

@Nnesha @mathstudent55 @e.mccormick @Loser66 @ash2326

OpenStudy (loser66):

Do you know partial derivative? implicit derivative?

OpenStudy (anonymous):

sort of, we just started

OpenStudy (loser66):

the first one x^3 +y^3 = 1 if we solve y w.r.t.x, it is hard to take derivative. Hence, we apply implicit one derivative both sides 3x^2 + 3y^2y' =0 that gives us y'= -x^2/y right?

OpenStudy (loser66):

that is it.

OpenStudy (anonymous):

oh, cool

OpenStudy (anonymous):

would it be positive?

OpenStudy (anonymous):

oh wait, nvm

OpenStudy (loser66):

for the second one x +cos (x+y) =0 same 1-sin(x+y) *(x+y)' =0 1+ sin(x+y)(1+y') =0 so, just solve for y' that is it. do the same with the rest/

OpenStudy (loser66):

oh, I mess up at the second line, it is 1- sin...

OpenStudy (anonymous):

Oh, gotcha. Is it okay if you help me with another question? If x=e^theta cos theta and y=e^theta sin theta, then, when theta=pi/2, dy/dx is...